OpenStudy (anonymous):

Find the distance between the points given. (2, 2) and (5, 5) 3 3√2 7√2

OpenStudy (anonymous):

distance formula attached

OpenStudy (anonymous):

\[\sqrt{(5-2)^2+(5-2)^2}\] \[\sqrt{(3)^2+(3)^2}\] \[\sqrt{18}\] d = \[3\sqrt{2}\] Are you understanding these? or are you just checking homework or something

OpenStudy (anonymous):

yes, and I understand them just some of them im not so sure about and im double cheeking if i did the work right,

OpenStudy (anonymous):

Alright good, practice makes perfect!

OpenStudy (anonymous):

Thats right !!:)